OpenStudy (anonymous):

What is the solution to the following system of equations? y = −x2 − 5x − 4 y = −x2 + 9x − 18 (−1, −10) (1, −10) (−1, 10) (1, 10) What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

OpenStudy (timaashorty):

\[1. \ ( 1,-10)\] Plug them back into the equations: -10 = −1(2) − 5(1) − 4✔ -10 = -(1)^2 + 9(1) - 18✔

OpenStudy (anonymous):

what about the second one

OpenStudy (timaashorty):

\[2.\ (-7,-2) \] Plug them: (-2) 2 + (-7)2 = 53 (-2) − (-7) = 5